# Memory in neural activity: long-range order without criticality

### v4

@@ V Discussion
While similar phenomena are often attributed to self-organized criticality (SOC) [29], self-organized bistability (SOB) [30], quasi-criticality [31, 32, 30], or the Griffiths phase [33, 34], we emphasize that the observed memory-induced long-range order (LRO) in our system is distinct and cannot be fully captured within these existing frameworks.
@@ VI.1 Numerical Details
Eqs. 1 and 2 are numerically integrated using the Euler-Maruyama method with a fixed time step, $\Delta t=10^{-2}$. The Gaussian white noise is simulated as a Wiener process with zero mean and variance $\Delta t$ at each integration step. Initial conditions $\rho_{i}(0)$ and $R_{i}(0)$ were sampled from narrow Gaussian distributions centered at low values of $\rho_{i}$ and $R_{i}$ with variance 0.01. This choice minimizes the transient time while yielding the same characteristic (spiking) post-transient dynamics. Periodic boundary conditions are imposed at the edges of the lattice. Movies visualizing the dynamics for various values of $\tau_{D}$ are included in the SI. Lattice sizes $N\equiv L^{2}$ of $64^{2}$, $128^{2}$, $256^{2}$, and $512^{2}$ were simulated. We note that $\rho$ and $R$ are cut off at 0 from below, giving a positive bias to the otherwise Gaussian noise when $\rho$ and/or $R$ are small. For all simulations in the main text, we chose the parameters $a=1$, $b=1.5$, $c=1$, $h=10^{-7}$, $D=1$, $\sigma=0.1$, $\delta=0.004$, $\epsilon=0.5$, $\Delta t=0.01$, and $\Delta_{tw}=0.3$ (generally in agreement with previous work [21]) We study the effect of varying parameters $a$ and $\sigma$ in the SI.
@@ VI.2 Neuronal Avalanches
[…]
@@ Appendix A Generalized Coarse-Graining Approach
$\left\{\begin{aligned} \dot{\rho}_{i}(t)&=R_{i}\rho_{i}+D\nabla^{2}\rho_{i}\;,\\ \dot{R}_{i}(t)&=-\frac{1}{\tau_{D}}R_{i}\rho_{i}\;.\end{aligned}\right$. (1)
@@
$R_{i}(N\tau_{D})$ $=R_{i}\big{(}(N-1)\tau_{D}\big{)}+\int_{(N-1)\tau_{D}}^{N\tau_{D}}dt\dot{R}_{i}(t)$ (2)
$\approx R_{i}\big{(}(N-1)\tau_{D}\big{)}(1-\overline{\rho_{i,N-1}})$
$\approx R_{i}\big{(}(N-2)\tau_{D}\big{)}(1-\overline{\rho_{i,N-2}})(1-\overline{\rho_{i,N-1}})$
$\approx R_{i}(0)\prod_{l=0}^{N-1}(1-\overline{\rho_{i,l}})\;,$
$\rho_{i}(N$ $\tau_{D})$ (3)
$\approx\rho_{i}\big{(}(N-1)\tau_{D}\big{)}+\int_{(N-1)\tau_{D}}^{N\tau_{D}}dt\dot{\rho}_{i}(t)$
$\approx\rho_{i}\big{(}(N-1)\tau_{D}\big{)}$
$\quad+\tau_{D}\bigg{(}R_{i}\big{(}(N-1)\tau_{D}\big{)}\overline{\rho_{i,N-1}}+D\nabla^{2}\overline{\rho_{i,N-1}}\bigg{)}$
$\approx\rho_{i}\big{(}(N-1)\tau_{D}\big{)}$
$\quad+\tau_{D}\bigg{(}R_{i}(0)\overline{\rho_{i,N-1}}\prod_{l=0}^{N-2}(1-\overline{\rho_{i,l}})+D\nabla^{2}\overline{\rho_{i,N-1}}\bigg{)}$
$\approx\rho_{i}(0)$
$\quad+\tau_{D}\sum_{m=0}^{N-1}\bigg{(}R_{i}(0)\overline{\rho_{i,m}}\prod_{l=0}^{m-1}(1-\overline{\rho_{i,l}})+D\nabla^{2}\overline{\rho_{i,m}}\bigg{)}.$
For $m=0$, replace $\prod_{l=0}^{m-1}(1-\overline{\rho_{i,l}})$ with $1$. Above, we extrapolate the forms of $R_{i}(N\tau_{D})$ and $\rho_{i}(N\tau_{D})$ by determining their forms in terms of $R_{i}((N-1)\tau_{D})$ and $\rho_{i}((N-1)\tau_{D})$, and then iterating.
@@
$\tau_{D}R_{i}(0)\overline{\rho_{i,m}}\prod_{l=0}^{m-1}(-\overline{\rho_{i,l}})\,$, (4)
@@
Let us now more explicitly investigate the discrete average $\overline{\rho_{i,k}}$ that was introduced in the main text. First, we study $\rho_{i}(k\tau_{D}+p/D)$, the constituent terms in its summation. For the case of $k=0$,
$\rho_{i}(1/D)$ $=\rho_{i}(0)+\int_{0}^{1/D}dt\dot{\rho}_{i}(t)$ (5)
$\approx\rho_{i}(0)+\frac{1}{D}\bigg{(}R_{i}(0)\rho_{i}(0)+D\nabla^{2}\rho_{i}(0)\bigg{)}$
$=\rho_{i}(0)\big{(}1+R_{i}(0)/D\big{)}+\sum_{j\in n.n.i}\bigg{(}\rho_{j}(0)-\rho_{i}(0)\bigg{)}\,$,
$\rho_{i}(2/D)$ $=\rho_{i}(1/D)+\int_{1/D}^{2/D}dt\dot{\rho}_{i}(t)$
$\approx\rho_{i}(1/D)+\frac{1}{D}\bigg{(}R_{i}(0)\rho_{i}(1/D)+D\nabla^{2}\rho_{i}(1/D)\bigg{)}$
$=\cdots+\sum_{j\in n.n.i}\Bigg{[}\,\sum_{k\in n.n.j}\bigg{(}\rho_{k}(0)-\rho_{j}(0)\bigg{)}$
$\quad\quad\quad\quad-\sum_{j\in n.n.i}\bigg{(}\rho_{j}(0)-\rho_{i}(0)\bigg{)}\Bigg{]}\,.$
[…]
This further generalizes for $k\neq 0$ (times longer than one resource timescale). In those instances, the form of $\rho_{i}(k\tau_{D}+p/D)$ will be even more complicated, as the value of $R_{i}(t)$ will vary during the dynamics. However, diffusive coupling will still cause individual activities to implicitly depend on the state of other activities farther and farther away. In this case, the radius over which implicit dependencies exist is $nN=N\lfloor\tau_{D}D\rfloor$, which is proportional to the total number of resource timesteps that have passed and the ratio between the activity and resource timescales.
To clarify, it is not this exponentially weakening radius of interactions due to the diffusive coupling that we interpret as LRO. Let us return to Eq. 4. Now, it is clear that $\overline{\rho_{i,k}}$ depends on the state of other activities within a wide radius of influence, albeit weakly. So, despite the contribution of individual activity regions being exponentially weak, this product series term suggests true, long-range activity-activity coupling up to order $N$, as each $\overline{\rho_{i,k}}$ depends implicitly on activies far away. This makes sense, as it is the resource (memory) variables that induce this LRO, which “update” $N$ times over a duration of time $N\tau_{D}$.
@@ B.1 Phase Structure of Cortical Dynamics
Instead, noise and diffusion are crucial factors in the down phase. Considering Eqs. 1 and 2 on an $L\times L$ lattice with periodic boundary condition (PBC), we can derive the dynamics of the mean fields, $\overline{\rho}=\sum_{i}\rho_{i}/L^{2}$ and $\overline{R}=\sum_{i}R_{i}/L^{2}$, by averaging over the entire lattice:
@@
$\dot{\overline{R}}=$ $\delta-\frac{1}{\tau_{D}}(\overline{R\rho}+\frac{\sigma}{L}\eta)\,.$
@@
$\dot{\overline{R}}=$ $\delta-\frac{1}{\tau_{D}}(\overline{R}\overline{\rho}+\mathrm{cov}(\rho,R)+\frac{\sigma}{L}\eta)\,.$
Here, we introduce some approximations to simplify our understanding of the system. First, for large systems where $L\to\infty$, the term $\frac{\sigma}{L}\eta$ approaches 0. Additionally, numerical evidence from the down phase indicates that both $\rho_{i}$ and $R_{i}$ experience minimal fluctuations around their equilibrium values primarily due to the noise term $\sigma\eta_{i}$. Consequently, we approximate $\mathrm{cov}(\rho,R)\approx 0$, and $\sigma_{\rho}^{2}\gg c\gamma_{\rho}\sigma_{\rho}^{3}\approx 0$. Numerical simulations at $L=16$, $\tau_{D}=10$ provide values $\sigma_{\rho}=0.0353$, $\mathrm{cov}(\rho,R)=9.84\times 10^{-5}$ and $\gamma_{\rho}=0.410$, supporting the validity of these approximations. With these simplifications, the dynamics of the mean fields are described by
@@
Next, we analytically estimate the value of $\sigma_{\rho}$. Assuming that, in the down phase, the fluctuation of $\rho_{i}$ around its equilibrium follows a Gaussian distribution, $\Delta\rho_{i}\sim N(0,\sigma_{\rho}^{2})$. With a small fluctuation $\Delta\rho_{i}$, we proceed by linearizing the dynamics of a single neural activity region, Eq. 1, around its equilibrium point. At equilibrium, $(-a+R_{i})\rho_{i}+b\rho_{i}^{2}-c\rho_{i}^{3}+h=0$, and
@@
$\dot{\rho}_{i}=$ $(-a+R_{i})\Delta\rho_{i}+2b\rho_{i}\Delta\rho_{i}-3c\rho_{i}^{2}\Delta\rho_{i}$ (11)
$+D\nabla^{2}(\Delta\rho_{i})+\sigma\eta_{i}\,.$
@@
$\Delta\rho_{i}(t+\Delta t)$ $=\Delta\rho_{i}(t)+\Big{(}(R_{i}-a+2b\rho_{i}-3c\rho_{i}^{2})\Delta\rho_{i}(t)$ (12)
$+D$ $\sum_{\langle i,j\rangle}(\Delta\rho_{j}(t)-\Delta\rho_{i}(t))\Big{)}\Delta t+\sigma\eta_{i}\sqrt{\Delta t}\,$,
$=\Delta\rho_{i}(t)(1-(4D-\alpha_{i})\Delta t)$
$\qquad+D\Delta t\sum_{\langle i,j\rangle}\Delta\rho_{j}(t)+\sigma\eta_{i}\sqrt{\Delta t}\,$,
where $\alpha_{i}=R_{i}-a+2b\rho_{i}-3c\rho_{i}^{2}$.
Assuming that $\Delta\rho_{i}$ and $\Delta\rho_{j}$ are uncorrelated and utilizing $\Delta\rho_{i}\sim N(0,\sigma_{\rho}^{2})$, we get:
$N(0,\sigma_{\rho}^{2})=$ $N(0,((1-(4D-\alpha_{i})\Delta t)\sigma_{\rho})^{2})$ (13)
@@
$\sigma_{\rho}^{2}=(1-(4D-\alpha_{i})\Delta t)^{2}\sigma_{\rho}^{2}+4D^{2}\Delta t^{2}\sigma_{\rho}^{2}+\sigma^{2}\Delta t\,$, (14)
$\implies$ $\sigma_{\rho}^{2}=\frac{\sigma^{2}}{8D-2\alpha_{i}-((4D-\alpha_{i})^{2}+4D^{2})\Delta t}\,.$
Note that $\alpha_{i}$ determines the stability of the equilibrium point in the single-region dynamics: $\alpha_{i}>0$ indicates an unstable equilibrium, while $\alpha_{i}<0$ denotes stability. Near the transition between the LRO phase and the down phase, $\alpha_{i}$ is approximately 0, a finding supported by numerical results. In addition, with $\Delta t\ll D$, the variance $\sigma_{\rho}^{2}$ simplifies to
@@
Thus, we observe a down phase with no neural activities below $\tau_{D}=8.15$, while the lower bound noted in numerical experiments is $\tau_{D}\approx 25$. This discrepancy could be attributed to several factors: for instance, minor $\rho$ activities may fail to reach the predefined threshold $\rho_{\mathrm{thrs}}=0.5$ and go unrecorded in the numerical experiments; additionally, some assumptions, such as $\Delta\rho_{i}\sim N(0,\sigma_{\rho}^{2})$, may not hold at slightly larger $\tau_{D}$ values.

### v5

@@ V Discussion
A natural follow-up question is the following: what is the role that either the dimensionality or structure of the lattice plays in inducing LRO via memory in such a system?
We expect that the presence of such memory-induced LRO should generalize to higher dimensions as well as some alternative lattice structures, since our analytical understanding of this phenomenon does not explicitly depend on the lattice connectivity itself (cf. Sec. III). On the contrary, this effect depends on (i) the existence of sufficiently strong memory (i.e., that the effective timescale of the memory is long) and (ii) that a sufficient number of pathways in the lattice exist whereby the memory effect can propagate through, correlating distant lattice sites. With (ii) in mind, we anticipate that the coordination number (number of nearest-neighbors per site) is the relevant quantity which must be sufficiently large to give enough “pathways” for the memory effect to propagate. This, of course, depends on the lattice structure/dimension.
For example, since the coordination number is 6 for a 3D square lattice as opposed to 4 for the 2D case, we expect it to be easier for the memory-induced LRO state to form in a 3D lattice than a 2D one. On the other hand, LRO should not emerge in the 1D case. There, only a single path exists through which the memory effect can correlate any pair of sites. So, whenever the correlations become significant, all sites should become correlated, inducing a rigid state which is characterized by exclusively system-wide avalanches. Thus, there appears to be no way to obtain the power-law distribution of avalanches which characterizes LRO in a 1D lattice.
Additionally, we must emphasize that while similar phenomena are often attributed to self-organized criticality (SOC) [29], self-organized bistability (SOB) [30], quasi-criticality [31, 32, 30], or the Griffiths phase [33, 34], the observed memory-induced long-range order (LRO) in our system is distinct and cannot be fully captured within these existing frameworks.
@@ VI.1 Numerical Details
Eqs. 1 and 2 are numerically integrated using the Euler-Maruyama method with a fixed time step, $\Delta t=10^{-2}$. The Gaussian white noise is simulated as a Wiener process with zero mean and variance $\Delta t$ at each integration step. Initial conditions $\rho_{\vec{x}}(0)$ and $R_{\vec{x}}(0)$ were sampled from narrow Gaussian distributions centered at low values of $\rho_{\vec{x}}$ and $R_{\vec{x}}$ with variance 0.01. This choice minimizes the transient time while yielding the same characteristic (spiking) post-transient dynamics. Periodic boundary conditions are imposed at the edges of the lattice. Movies visualizing the dynamics for various values of $\tau_{D}$ are included in the SI. Lattice sizes $N\equiv L^{2}$ of $64^{2}$, $128^{2}$, $256^{2}$, and $512^{2}$ were simulated. We note that $\rho$ and $R$ are cut off at 0 from below, giving a positive bias to the otherwise Gaussian noise when $\rho$ and/or $R$ are small. For all simulations in the main text, we chose the parameters $a=1$, $b=1.5$, $c=1$, $h=10^{-7}$, $D=1$, $\sigma=0.1$, $\delta=0.004$, $\epsilon=0.5$, $\Delta t=0.01$, and $\Delta_{tw}=0.3$ (generally in agreement with previous work [21]) We study the effect of varying parameters $a$ and $\sigma$ in the SI.
@@ VI.2 Neuronal Avalanches
In our simulated lattice of neural activity, we define avalanches as follows. When the activity at region $\vec{x}$, $\rho_{\vec{x}}(t)$, crosses a threshold $\epsilon$, we say an event has occurred at time $t$ and location $\vec{x}$. We then search for similar activity-crossing events at the four nearest-neighbors of $\vec{x}$ in the lattice during the following time $\Delta_{tw}$, called the “time window”. Within the whole time window (from t to t + $\Delta_{tw}$), if an event is found at time $t^{\prime}$, the size of the avalanche is increased by 1, and further events are searched for adjacent to the newest event between times $t^{\prime}$ to $t^{\prime}+\Delta_{tw}$. An avalanche stops when no events are found in any of the nearest-neighbors of recent (within a time $\Delta_{tw}$) events in that avalanche. As a result, a lone activity spike would characterize an avalanche of size 1.
@@ Appendix A Generalized Coarse-Graining Approach
$\left\{\begin{aligned} \dot{\rho}_{\vec{x}}(t)&=R_{\vec{x}}\rho_{\vec{x}}+D\nabla^{2}\rho_{\vec{x}}\;,\\ \dot{R}_{\vec{x}}(t)&=-\frac{1}{\tau_{D}}R_{\vec{x}}\rho_{\vec{x}}\;.\end{aligned}\right$. (1)
@@
$R_{\vec{x}}(N\tau_{D})$ $=R_{\vec{x}}\big((N-1)\tau_{D}\big)+\int_{(N-1)\tau_{D}}^{N\tau_{D}}dt\dot{R}_{\vec{x}}(t)$ (2)
$\approx R_{\vec{x}}\big((N-1)\tau_{D}\big)(1-\overline{\rho_{\vec{x},N-1}})$
$\approx R_{\vec{x}}\big((N-2)\tau_{D}\big)(1-\overline{\rho_{\vec{x},N-2}})(1-\overline{\rho_{\vec{x},N-1}})$
$\approx R_{\vec{x}}(0)\prod_{l=0}^{N-1}(1-\overline{\rho_{\vec{x},l}})\;,$
$\rho_{\vec{x}}(N$ $\tau_{D})$ (3)
$\approx\rho_{\vec{x}}\big((N-1)\tau_{D}\big)+\int_{(N-1)\tau_{D}}^{N\tau_{D}}dt\dot{\rho}_{\vec{x}}(t)$
$\approx\rho_{\vec{x}}\big((N-1)\tau_{D}\big)$
$\quad+\tau_{D}\bigg(R_{\vec{x}}\big((N-1)\tau_{D}\big)\overline{\rho_{\vec{x},N-1}}+D\nabla^{2}\overline{\rho_{\vec{x},N-1}}\bigg)$
$\approx\rho_{\vec{x}}\big((N-1)\tau_{D}\big)$
$\quad+\tau_{D}\bigg(R_{\vec{x}}(0)\overline{\rho_{\vec{x},N-1}}\prod_{l=0}^{N-2}(1-\overline{\rho_{\vec{x},l}})+D\nabla^{2}\overline{\rho_{\vec{x},N-1}}\bigg)$
$\approx\rho_{\vec{x}}(0)$
$\quad+\tau_{D}\sum_{m=0}^{N-1}\bigg(R_{\vec{x}}(0)\overline{\rho_{\vec{x},m}}\prod_{l=0}^{m-1}(1-\overline{\rho_{\vec{x},l}})+D\nabla^{2}\overline{\rho_{\vec{x},m}}\bigg).$
For $m=0$, replace $\prod_{l=0}^{m-1}(1-\overline{\rho_{\vec{x},l}})$ with $1$. Above, we extrapolate the forms of $R_{\vec{x}}(N\tau_{D})$ and $\rho_{\vec{x}}(N\tau_{D})$ by determining their forms in terms of $R_{\vec{x}}((N-1)\tau_{D})$ and $\rho_{\vec{x}}((N-1)\tau_{D})$, and then iterating.
@@
$\tau_{D}R_{\vec{x}}(0)\overline{\rho_{\vec{x},m}}\prod_{l=0}^{m-1}(-\overline{\rho_{\vec{x},l}})\,$, (4)
@@
Let us now more explicitly investigate the discrete average $\overline{\rho_{\vec{x},k}}$ that was introduced in the main text. First, we study $\rho_{\vec{x}}(k\tau_{D}+p/D)$, the constituent terms in its summation. For the case of $k=0$,
$\rho_{\vec{x}}(1/D)$ $=\rho_{\vec{x}}(0)+\int_{0}^{1/D}dt\dot{\rho}_{\vec{x}}(t)$ (5)
$\approx\rho_{\vec{x}}(0)+\frac{1}{D}\bigg(R_{\vec{x}}(0)\rho_{\vec{x}}(0)+D\nabla^{2}\rho_{\vec{x}}(0)\bigg)$
$=\rho_{\vec{x}}(0)\big(1+R_{\vec{x}}(0)/D\big)+\sum_{\vec{y}\in n.n.\vec{x}}\bigg(\rho_{\vec{y}}(0)-\rho_{\vec{x}}(0)\bigg)\,$,
$\rho_{\vec{x}}(2/D)$ $=\rho_{\vec{x}}(1/D)+\int_{1/D}^{2/D}dt\dot{\rho}_{\vec{x}}(t)$
$\approx\rho_{\vec{x}}(1/D)+\frac{1}{D}\bigg(R_{\vec{x}}(0)\rho_{\vec{x}}(1/D)+D\nabla^{2}\rho_{\vec{x}}(1/D)\bigg)$
$=\cdots+\sum_{\vec{y}\in n.n.\vec{x}}\Bigg[\,\sum_{\vec{z}\in n.n.\vec{y}}\bigg(\rho_{\vec{z}}(0)-\rho_{\vec{y}}(0)\bigg)$
$\quad\quad\quad\quad-\sum_{\vec{y}\in n.n.\vec{x}}\bigg(\rho_{\vec{y}}(0)-\rho_{\vec{x}}(0)\bigg)\Bigg]\,.$
Therefore, after 2 activity timesteps, because each neighbor at site $\vec{x}$ will have had time to update due to local interactions with its neighbors, the activity at site $\vec{x}$ will depend on all the neighbors of the activities at $\vec{y}$ as well (denoted here with $\vec{z}$). This pattern will continue so that over a timescale $\tau_{D}$, each activity will implicitly depend on all its neighbors within $n$ lattice sites away (again, $n\equiv\lfloor\tau_{D}D\rfloor$ is the ratio of the resource to activity timescales). This occurs at every point in the lattice simultaneously.
This further generalizes for $k\neq 0$ (times longer than one resource timescale). In those instances, the form of $\rho_{\vec{x}}(k\tau_{D}+p/D)$ will be even more complicated, as the value of $R_{\vec{x}}(t)$ will vary during the dynamics. However, diffusive coupling will still cause individual activities to implicitly depend on the state of other activities farther and farther away. In this case, the radius over which implicit dependencies exist is $nN=N\lfloor\tau_{D}D\rfloor$, which is proportional to the total number of resource timesteps that have passed and the ratio between the activity and resource timescales.
To clarify, it is not this exponentially weakening radius of interactions due to the diffusive coupling that we interpret as LRO. Let us return to Eq. 4. Now, it is clear that $\overline{\rho_{\vec{x},k}}$ depends on the state of other activities within a wide radius of influence, albeit weakly. So, despite the contribution of individual activity regions being exponentially weak, this product series term suggests true, long-range activity-activity coupling up to order $N$, as each $\overline{\rho_{\vec{x},k}}$ depends implicitly on activies far away. This makes sense, as it is the resource (memory) variables that induce this LRO, which “update” $N$ times over a duration of time $N\tau_{D}$.
@@ B.1 Phase Structure of Cortical Dynamics
Instead, noise and diffusion are crucial factors in the down phase. Considering Eqs. 1 and 2 on an $L\times L$ lattice with periodic boundary condition (PBC), we can derive the dynamics of the mean fields, $\overline{\rho}=\sum_{\vec{x}}\rho_{\vec{x}}/L^{2}$ and $\overline{R}=\sum_{\vec{x}}R_{\vec{x}}/L^{2}$, by averaging over the entire lattice:
@@
$\dot{\overline{R}}=$ $\delta-\frac{1}{\tau_{D}}(\overline{R\rho}+\frac{\sigma}{L}\zeta)\,.$
@@
$\dot{\overline{R}}=$ $\delta-\frac{1}{\tau_{D}}(\overline{R}\overline{\rho}+\mathrm{cov}(\rho,R)+\frac{\sigma}{L}\zeta)\,.$
Here, we introduce some approximations to simplify our understanding of the system. First, for large systems where $L\to\infty$, the terms $\frac{\sigma}{L}\eta$ and $\frac{\sigma}{L}\zeta$ approach 0. Additionally, numerical evidence from the down phase indicates that both $\rho_{\vec{x}}$ and $R_{\vec{x}}$ experience minimal fluctuations around their equilibrium values primarily due to the noise terms $\sigma\eta_{\vec{x}}$ and $\sigma\zeta_{\vec{x}}$. Consequently, we approximate $\mathrm{cov}(\rho,R)\approx 0$, and $\sigma_{\rho}^{2}\gg c\gamma_{\rho}\sigma_{\rho}^{3}\approx 0$. Numerical simulations at $L=16$, $\tau_{D}=10$ provide values $\sigma_{\rho}=0.0353$, $\mathrm{cov}(\rho,R)=9.84\times 10^{-5}$ and $\gamma_{\rho}=0.410$, supporting the validity of these approximations. With these simplifications, the dynamics of the mean fields are described by
@@
Next, we analytically estimate the value of $\sigma_{\rho}$. Assuming that, in the down phase, the fluctuation of $\rho_{\vec{x}}$ around its equilibrium follows a Gaussian distribution, $\Delta\rho_{\vec{x}}\sim N(0,\sigma_{\rho}^{2})$. With a small fluctuation $\Delta\rho_{\vec{x}}$, we proceed by linearizing the dynamics of a single neural activity region, Eq. 1, around its equilibrium point. At equilibrium, $(-a+R_{\vec{x}})\rho_{\vec{x}}+b\rho_{\vec{x}}^{2}-c\rho_{\vec{x}}^{3}+h=0$, and
@@
$\dot{\rho}_{\vec{x}}=$ $(-a+R_{\vec{x}})\Delta\rho_{\vec{x}}+2b\rho_{\vec{x}}\Delta\rho_{\vec{x}}-3c\rho_{\vec{x}}^{2}\Delta\rho_{\vec{x}}$ (11)
$+D\nabla^{2}(\Delta\rho_{\vec{x}})+\sigma\eta_{\vec{x}}\,.$
@@
$\Delta\rho_{\vec{x}}(t+\Delta t)$ $=\Delta\rho_{\vec{x}}(t)+\Big((R_{\vec{x}}-a+2b\rho_{\vec{x}}-3c\rho_{\vec{x}}^{2})\Delta\rho_{\vec{x}}(t)$ (12)
$+D$ $\sum_{\vec{y}\in n.n.\vec{x}}(\Delta\rho_{\vec{y}}(t)-\Delta\rho_{\vec{x}}(t))\Big)\Delta t+\sigma\eta_{\vec{x}}\sqrt{\Delta t}\,$,
$=\Delta\rho_{\vec{x}}(t)(1-(4D-\alpha_{\vec{x}})\Delta t)$
$\qquad+D\Delta t\sum_{\vec{y}\in n.n.\vec{x}}\Delta\rho_{\vec{y}}(t)+\sigma\eta_{\vec{x}}\sqrt{\Delta t}\,$,
where $\alpha_{\vec{x}}=R_{\vec{x}}-a+2b\rho_{\vec{x}}-3c\rho_{\vec{x}}^{2}$.
Assuming that $\Delta\rho_{\vec{x}}$ and $\Delta\rho_{\vec{y}}$ are uncorrelated and utilizing $\Delta\rho_{\vec{x}}\sim N(0,\sigma_{\rho}^{2})$, we get:
$N(0,\sigma_{\rho}^{2})=$ $N(0,((1-(4D-\alpha_{\vec{x}})\Delta t)\sigma_{\rho})^{2})$ (13)
@@
$\sigma_{\rho}^{2}=(1-(4D-\alpha_{\vec{x}})\Delta t)^{2}\sigma_{\rho}^{2}+4D^{2}\Delta t^{2}\sigma_{\rho}^{2}+\sigma^{2}\Delta t\,$, (14)
$\implies$ $\sigma_{\rho}^{2}=\frac{\sigma^{2}}{8D-2\alpha_{\vec{x}}-((4D-\alpha_{\vec{x}})^{2}+4D^{2})\Delta t}\,.$
Note that $\alpha_{\vec{x}}$ determines the stability of the equilibrium point in the single-region dynamics: $\alpha_{\vec{x}}>0$ indicates an unstable equilibrium, while $\alpha_{\vec{x}}<0$ denotes stability. Near the transition between the LRO phase and the down phase, $\alpha_{\vec{x}}$ is approximately 0, a finding supported by numerical results. In addition, with $\Delta t\ll D$, the variance $\sigma_{\rho}^{2}$ simplifies to
@@
Thus, we observe a down phase with no neural activities below $\tau_{D}=8.15$, while the lower bound noted in numerical experiments is $\tau_{D}\approx 25$. This discrepancy could be attributed to several factors: for instance, minor $\rho$ activities may fail to reach the predefined threshold $\rho_{\mathrm{thrs}}=0.5$ and go unrecorded in the numerical experiments; additionally, some assumptions, such as $\Delta\rho_{\vec{x}}\sim N(0,\sigma_{\rho}^{2})$, may not hold at slightly larger $\tau_{D}$ values.